# On the singular loci of higher secant varieties of Veronese embeddings

### v2

@@ 2.3. Estimate for the linear span of tangents moving along a subsecant variety
First, we give an explicit description of the embedded tangent space $\mathbb{T}_{x}v_{d}({\mathbb{P}}^{n})\subset{\mathbb{P}}^{N}$ at a point $x$ in $v_{d}({\mathbb{P}}^{n})$ or $v_{d}({\mathbb{P}}^{m})$ as follows (see also [14]).
Recall that $\mathsf{mono}[t]_{\leq e}$ denotes the set of monomials $f\in{\mathbb{C}}[t_{1},\dots,t_{m}]$ with $\deg f\leq e$. Note that $1\in\mathsf{mono}[t]_{\leq e}$ as the monomial of degree 00. As mentioned in Remark 8, as changing homogeneous coordinates $t_{0},t_{1},\dots,t_{m},u_{1},u_{2},\dots,u_{m^{\prime}}$ on ${\mathbb{P}}^{n}$ with $m^{\prime}:=n-m$, we may assume that ${\mathbb{P}}^{m}$ is the zero set of $u_{1},\dots,u_{m^{\prime}}$. On the affine open subset $\left\{t_{0}\neq 0\right\}$, the Veronese embedding $v_{d}:{\mathbb{P}}^{n}\rightarrow{\mathbb{P}}^{N}$ is parameterized by monomials of ${\mathbb{C}}[t_{1},\dots,t_{m},u_{1},\dots,u_{m^{\prime}}]$ of degree $\leq d$. So it is expressed as
(11) $[\,\mathsf{mono}[t]_{\leq d}:u_{1}\cdot\mathsf{mono}[t]_{\leq d-1}:\quad\cdots\quad:u_{m^{\prime}}\cdot\mathsf{mono}[t]_{\leq d-1}:\quad*\quad\,]~{},$
where $u_{i}\cdot\mathsf{mono}[t]_{\leq d-1}$ means $\left\{u_{i}f\mid f\in\mathsf{mono}[t]_{\leq d-1}\right\}=(u_{i}:u_{i}t_{1}:u_{i}t_{2}:\dots:u_{i}t_{m}^{d-1})$, and “*” means the remaining monomials.
@@
$\begin{bmatrix}v_{d}\\ \partial v_{d}/\partial t_{1}\\ \vdots\\ \partial v_{d}/\partial t_{m}\\ \partial v_{d}/\partial u_{1}\\ \vdots\\ \partial v_{d}/\partial u_{m^{\prime}}\end{bmatrix}(x^{\prime})$ =[$\mathsf{mono}[t]_{\leq d}&u_{1}\cdot\mathsf{mono}[t]_{\leq d-1}&\dots&u_{m^{\prime}}\cdot\mathsf{mono}[t]_{\leq d-1}&&*&\\ (\mathsf{mono}[t]_{\leq d})_{t_{1}}&u_{1}\cdot(\mathsf{mono}[t]_{\leq d-1})_{t_{1}}&\dots&u_{m^{\prime}}\cdot(\mathsf{mono}[t]_{\leq d-1})_{t_{1}}&&*&\\ \vdots&\vdots&&\vdots&&\vdots&\\ (\mathsf{mono}[t]_{\leq d})_{t_{m}}&u_{1}\cdot(\mathsf{mono}[t]_{\leq d-1})_{t_{m}}&\dots&u_{m^{\prime}}\cdot(\mathsf{mono}[t]_{\leq d-1})_{t_{m}}&&*&\\ \mathsf{O}&\mathsf{mono}[t]_{\leq d-1}&\dots&\mathsf{O}&&*&\\ \vdots&&\ddots&&&\vdots&\\ \mathsf{O}&\mathsf{O}&\dots&\mathsf{mono}[t]_{\leq d-1}&&*&\end{array}\right](x^{\prime})~{}$
@@
As a consequence, we settle a key proposition which estimates a lower bound of dimension of linear span of moving embedded tangent spaces along a subset of a given ${\mathbb{P}}^{m}$.
@@
$B_{0}=v_{d}(A),B_{1}=(\partial v_{d}/\partial u_{1})(A),~{}\dots~{},B_{m^{\prime}}=(\partial v_{d}/\partial u_{m^{\prime}})(A)$
@@
(15) $\langle\,\Lambda\cup B_{0},B_{1},~{}\dots~{},B_{m^{\prime}}\,\rangle\subset{\mathbb{P}}^{N}$
is of dimension $\dim(\langle\,\Lambda\cup B_{0}\,\rangle)+\dim(\langle\,B_{1}\,\rangle)+\dots\dim(\langle\,B_{m^{\prime}}\,\rangle)+m^{\prime}$.
By (13) again, we see that $B_{0}\simeq v_{d}(A)$ and $B_{i}\simeq v_{d-1,m}(A)$ for $1\leq i\leq m^{\prime}$. Since the linear variety (15) is contained in $\langle\,\Lambda\cup\bigcup_{x\in v_{d}(A)}\mathbb{T}_{x}(v_{d}({\mathbb{P}}^{n}))\,\rangle$, we have the assertion. ∎
@@ 3.1. Symmetric flattening and conormal space computation
For the proof of Theorem 1, we begin with some preliminaries on equations for secant varieties and conormal space computation via known sets of equations, whereas we mainly adopt geometric viewpoint and techniques for $m\geq 2$ case in §4.
Let $V$ be a $(n+1)$-dimensional ${\mathbb{C}}$-vector space ${\mathbb{C}}\langle x_{0},x_{1},\dots,x_{n}\rangle$. For the Veronese variety $v_{d}({\mathbb{P}}V)$, we have a natural 1-1 correspondence between a point of the ambient space $\langle\,v_{d}({\mathbb{P}}V)\,\rangle$ and a homogeneous polynomial $f$ of degree $d$ (or a form) in $S:={\mathbb{C}}[x_{0},x_{1},\dots,x_{n}]$. First of all, let us recall some related notions to this correspondence.
Given a form $f$ of degree $d$, the minimum number of linear forms $l_{i}$ needed to write $f$ as a sum of $d$-th powers is the so-called (Waring) rank of $f$ and let us denote it by $\operatorname{rank}(f)$. The (Waring) border rank is given by the same notion in the limiting sense. In other words, if there is a family $\{f_{\epsilon}\mid\epsilon>0\}$ of polynomials with constant rank $r$ and $\lim_{\epsilon\to 0}f_{\epsilon}=f$, then we say that $f$ has border rank at most $r$. The minimum such $r$ is called the border rank of $f$ and let’s denote it again by $\underline{\mathrm{rank}}(f)$. Note that by definition $\sigma_{k}(v_{d}({\mathbb{P}}V))$ is the variety of homogeneous polynomials $f$ of degree $d$ with border rank $\underline{\mathrm{rank}}(f)\leq k$.
@@
$g\cdot f=g(\partial_{0},\partial_{1},\ldots,\partial_{n})f\in S_{a}~{},$
@@
$\mathbf{X}^{I}=\frac{1}{i_{0}!\cdots i_{n}!}x_{0}^{i_{0}}\cdots x_{n}^{i_{n}}$ $\quad\textrm{and}\quad\mathbf{Y}^{J}=y_{0}^{j_{0}}\cdots y_{n}^{j_{n}}~{},$
@@
($\ &b_{I,J}&\\ &&\end{array}\right)\quad\textrm{with $b_{I,J}=b_{I+J}$}~{},$
in the bases defined above. We call this ‘$(d-a,a)$-symmetric flattening (or catalecticant) matrix’ of $f$. It is easy to see that the transpose $\phi_{d-a,a}(f)^{T}$ is equal to $\phi_{a,d-a}(f)$.
It is obvious that if $f$ has (border) rank 1, then any symmetric flattening $\phi_{d-a,a}(f)$ has rank 1. By subadditivity of matrix rank, we also know that $\operatorname{rank}~{}\phi_{d-a,a}(f)\leq r$ if $\underline{\mathrm{rank}}(f)\leq r$. So, we obtain a set of defining equations coming from $(k+1)$-minors of the matrix $\phi_{d-a,a}(f)$ for $\sigma_{k}(v_{d}({\mathbb{P}}V))$. For some small values of $k$, it is known that these minors are sufficient to cut the variety $\sigma_{k}(v_{d}({\mathbb{P}}V))$ scheme-theoretically (see [23, theorem 3.2.1 (1)]).
Let us recall another basic terms and facts. Let $Z\subset{\mathbb{P}}W$ be a (reduced and irreducible) variety and $\hat{Z}$ be its affine cone in $W$. Consider a (closed) point $p\in\hat{Z}$ and say $[p]$ the corresponding point in ${\mathbb{P}}W$. We denote the affine tangent space to $Z$ at $[p]$ in $W$ by $\hat{T}_{[p]}Z$ and we define the (affine) conormal space to $Z$ at $[p]$, $\hat{N}^{\ast}_{[p]}Z$ as the annihilator $(\hat{T}_{[p]}Z)^{\perp}\subset W^{\ast}$. Since $\dim\hat{N}^{\ast}_{[p]}Z+\dim\hat{T}_{[p]}Z=\dim W$ and $\dim Z\leq\dim\hat{T}_{[p]}Z-1$, we get that $\dim\hat{N}^{\ast}_{[p]}Z\leq\operatorname{codim}(Z,{\mathbb{P}}W)$ and the equality holds if and only if $Z$ is smooth at $[p]$. This conormal space is quite useful to study the (embedded) tangent space $\mathbb{T}_{[p]}Z$.
For any given form $f\in S^{d}V$, we call $\partial\in T_{t}$ apolar to $f$ if the differentiation $\partial\cdot f$ gives zero (i.e. $\partial\in\ker\phi_{t,d-t}(f)$). And we define the apolar ideal $f^{\perp}\subset T$ as
$f^{\perp}:=\{\partial\in T~{}|~{}\partial\cdot f=0\}~{}.$
It is straightforward to see that $f^{\perp}$ is indeed an ideal of $T$. Moreover, it is well-known that the quotient ring $T_{f}:=T/f^{\perp}$ is an Artinian Gorenstein algebra with socle degree $d$ (see e.g. [19, chapter 1]). In terms of this apolar ideal, we have a useful description of (a part of) conormal space as follows:
@@
Suppose that $f\in S^{d}V$ corresponds to a (closed) point of $\sigma_{k}(v_{d}({\mathbb{P}}V))\setminus\sigma_{k-1}(v_{d}({\mathbb{P}}V))$. Then, for any $a$ with $1\leq a\leq{\lfloor\frac{d+1}{2}\rfloor}$ with $\operatorname{rank}\phi_{d-a,a}(f)=k$ we have
@@
Let $X\subset{\mathbb{P}}W$ be any variety. For any linear embedding $W\hookrightarrow A\otimes B$ and the induced embedding $X\subset{\mathbb{P}}W\hookrightarrow{\mathbb{P}}(A\otimes B)$, it is well-known that for any $f\in\sigma_{k}(X)\subset{\mathbb{P}}(A\otimes B)$, considering $\sigma_{k}(X)$ as a subvariety of ${\mathbb{P}}(A\otimes B)$, we have
$\hat{N}^{\ast}_{[f]}\sigma_{k}(X)\supseteq\ker(f)\otimes\operatorname{im}(f)^{\perp}=\hat{N}^{\ast}_{[f]}\sigma_{p}(Seg({\mathbb{P}}A\times{\mathbb{P}}B))\quad\textrm{in $A^{\ast}\otimes B^{\ast}$}$
provided that $X\subseteq\sigma_{p}(Seg({\mathbb{P}}A\times{\mathbb{P}}B))$, $X\nsubseteq\sigma_{p-1}(Seg({\mathbb{P}}A\times{\mathbb{P}}B))$ and $f$ has rank $k\cdot p$ as a linear map in $\operatorname{Hom}(A^{\ast},B)$ (see e.g. [23, section 2.5]). Here, $Seg({\mathbb{P}}A\times{\mathbb{P}}B)$ means the Segre variety in ${\mathbb{P}}(A\otimes B)$.
@@
The assertion is immediate when we apply this fact to a partial polarization $S^{d}V\hookrightarrow S^{a}V\otimes S^{d-a}V$, because $X=v_{d}({\mathbb{P}}V)$ is contained in $Seg({\mathbb{P}}S^{a}V\times{\mathbb{P}}S^{d-a}V)$ in ${\mathbb{P}}(S^{a}V\otimes S^{d-a}V)$ (i.e. $p=1$ case) and $\operatorname{rank}\phi_{d-a,a}(f)=k$, $\ker\phi_{d-a,a}(f)=(f^{\perp})_{d-a}$, $\operatorname{im}(\phi_{d-a,a}(f))^{\perp}=(f^{\perp})_{a}$. ∎
@@ 3.2. Proof of Theorem 1
(i) Let $f$ be any form belonging to $\sigma_{k}(v_{d}({\mathbb{P}}^{1}))\setminus\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$. Set $X=v_{d}({\mathbb{P}}^{n})\subset{\mathbb{P}}^{N}$, the Veronese variety. Consider $f$ as a polynomial in ${\mathbb{C}}[x_{0},x_{1}]$ as in §3.1. Then, by [19, theorem 1.44] we know that $T/f^{\perp}$ is an Artinian Gorenstein algebra with socle degree $d$ and that $f^{\perp}$ is a complete intersection of two homogeneous polynomials $F,G$ of each degree $a$ and $b$ $(a\leq b)$ with $a+b=d+2$ as an ideal of ${\mathbb{C}}[y_{0},y_{1}]$, where the Hilbert function of $T/f^{\perp}$ is $(1,2,\cdots,a-1,a,\cdots,a,a-1,\cdots,2,1)$. We claim that $\operatorname{rank}~{}\phi_{k,d-k}(f)=k$ (i.e. $a=k$). If $a<k$, then $\operatorname{rank}~{}\phi_{t,d-t}(f)<k$ for all $t$. Since it is well-known that the $k$-minors of catalecticant $\phi_{t,d-t}$ for any $k-1\leq t\leq d-(k-1)$ gives the ideal of $\sigma_{k-1}(v_{d}({\mathbb{P}}^{1}))$, we get $f\in\sigma_{k-1}(v_{d}({\mathbb{P}}^{1}))\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ which is contradiction.
@@
Now, let us show that $\sigma_{k}(X)$ is smooth at $f$ by computing the dimension of conormal space. In general, we have
(16) (${n+d\choose d}-kn-k\geq\dim_{\mathbb{C}}\hat{N}^{\ast}_{[f]}\sigma_{k}(X)~{},$
where the left hand side is given by the expected codimension of the conormal space. By Proposition 23, we also have
(17) $\dim_{\mathbb{C}}\hat{N}^{\ast}_{[f]}\sigma_{k}(X)\geq\dim_{\mathbb{C}}(f^{\perp})_{k}\cdot(f^{\perp})_{d-k}~{}.$
@@
Since $k\leq\frac{d+1}{2}$ by the assumption, note that $d-k\geq k$ unless $d$ is odd and $k=\frac{d+1}{2}$, in which $d-k=\frac{d-1}{2}<k$.
a) $d$: odd and $k=\frac{d+1}{2}$ case : In this case, we have
@@
b) $d$: even or $d$: odd and $k<\frac{d+1}{2}$ case : In the second case, we have $k\leq d-k$ and
@@
(ii) First note that $\dim\sigma_{k}(v_{d}({\mathbb{P}}^{1}))=\min\left\{2k-1,d\right\}$ and the incidence $I$ has dimension $12k-$ 1. In the case $d\leq 2k-2$, each fiber of $p:I\rightarrow{\mathbb{P}}^{d}$ is of dimension $\geq 1$, and then $q_{i}(p^{-1}(a))=v_{d}({\mathbb{P}}^{1})$ for some $i$ with $1\leq i\leq k$ in the incidence (5) in §2.2.
Now, let $n\geq 3,k=3$ or $n\geq 2,k\geq 4$ and $d=2k-2$. Suppose $\sigma_{k}(v_{d}({\mathbb{P}}^{1}))\not\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$. Then a general point $a\in\sigma_{k}(v_{d}({\mathbb{P}}^{1}))={\mathbb{P}}^{d}$ is a smooth point of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$. Since $q_{i}(p^{-1}(a))=v_{d}({\mathbb{P}}^{1})$ for some $i$, it follows from Lemma 12 that
(18) $\langle\,\bigcup_{x\in v_{d}({\mathbb{P}}^{1})}\mathbb{T}_{x}(v_{d}({\mathbb{P}}^{n}))\,\rangle\subset\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{n}))).$
Taking $m=1$ and $A={\mathbb{P}}^{1}$ in Proposition 22, the number (14), a lower bound for dimension of left hand side of (18), is equal to $dn$. Thus we have
$(2k-2)n=dn\leq k(n+1)-1,$

### v3

@@ 2.3. Estimate for the linear span of tangents moving along a subsecant variety
First, we give the following explicit description of the embedded tangent space $\mathbb{T}_{x}v_{d}({\mathbb{P}}^{n})\subset{\mathbb{P}}^{N}$ to $v_{d}({\mathbb{P}}^{n})$ at a point $x$ in $v_{d}({\mathbb{P}}^{n})$ or $v_{d}({\mathbb{P}}^{m})$. Note that it is related to computations of Gauss maps (see [FI17]).
Recall that $\mathsf{mono}[t]_{\leq e}$ denotes the set of monomials $f\in{\mathbb{C}}[t_{1},\dots,t_{m}]$ with $\deg f\leq e$. Then $1\in\mathsf{mono}[t]_{\leq e}$ as the monomial of degree 00. As mentioned in Remark 8, as changing homogeneous coordinates $t_{0},t_{1},\dots,t_{m},u_{1},u_{2},\dots,u_{m^{\prime}}$ on ${\mathbb{P}}^{n}$ with $m^{\prime}=n-m$, we may assume that ${\mathbb{P}}^{m}$ is the zero set of $u_{1},\dots,u_{m^{\prime}}$. On the affine open subset $\left\{t_{0}\neq 0\right\}$, the Veronese embedding $v_{d}:{\mathbb{P}}^{n}\rightarrow{\mathbb{P}}^{N}$ is parameterized by monomials of ${\mathbb{C}}[t_{1},\dots,t_{m},u_{1},\dots,u_{m^{\prime}}]$ of degree $\leq d$. So it is expressed as
(16) $[\,\mathsf{mono}[t]_{\leq d}:u_{1}\cdot\mathsf{mono}[t]_{\leq d-1}:\quad\cdots\quad:u_{m^{\prime}}\cdot\mathsf{mono}[t]_{\leq d-1}:\quad*\quad\,]\,,$
where $u_{i}\cdot\mathsf{mono}[t]_{\leq d-1}$ means
$\left\{u_{i}f\mid f\in\mathsf{mono}[t]_{\leq d-1}\right\}=(u_{i}:u_{i}t_{1}:u_{i}t_{2}:\dots:u_{i}t_{m}^{d-1}),$
and “*” means the remaining monomials.
@@
[…]
@@
As a consequence, we settle a key proposition which estimates a lower bound of the dimension of the linear span of moving embedded tangent spaces along a subset of a given ${\mathbb{P}}^{m}$.
@@
$B_{0}=v_{d}(A),B_{1}=(\partial v_{d}/\partial u_{1})(A),\,\dots\,,B_{m^{\prime}}=(\partial v_{d}/\partial u_{m^{\prime}})(A)$
@@
(20) $\langle\,\Lambda\cup B_{0},B_{1},\,\dots\,,B_{m^{\prime}}\,\rangle\subset{\mathbb{P}}^{N}$
is of dimension $\dim(\langle\,\Lambda\cup B_{0}\,\rangle)+\dim(\langle\,B_{1}\,\rangle)+\dots+\dim(\langle\,B_{m^{\prime}}\,\rangle)+m^{\prime}$.
Again, by (18) we see that $B_{0}\simeq v_{d}(A)$ and $B_{i}\simeq v_{d-1,m}(A)$ for $1\leq i\leq m^{\prime}$. Since the linear variety (20) is contained in $\langle\,\Lambda\cup\bigcup_{x\in v_{d}(A)}\mathbb{T}_{x}(v_{d}({\mathbb{P}}^{n}))\,\rangle$, we have the assertion. ∎
@@ 3.1. Symmetric flattening and conormal space computation
For the proof of Theorem 1, we begin with some preliminaries on equations for secant varieties and conormal space computation via known sets of equations, whereas we mainly adopt the geometric viewpoint and techniques for the $m\geq 2$ case in §4.
Let $V$ be a $(n+1)$-dimensional ${\mathbb{C}}$-vector space ${\mathbb{C}}\langle x_{0},x_{1},\dots,x_{n}\rangle$. Let $f\in S^{d}V$ be a homogeneous polynomial of degree $d$ (or $d$-form) and $[f]$ be the corresponding point in ${\mathbb{P}}S^{d}V$. In this paper, we frequently abuse notations as denoting both a $d$-form in $S^{d}V$ and the point in ${\mathbb{P}}S^{d}V$ just by $f$. For the Veronese variety $v_{d}({\mathbb{P}}V)$, we have a natural 1-1 correspondence between points of the ambient space $\langle\,v_{d}({\mathbb{P}}V)\,\rangle$ and equivalent classes of degree $d$-forms in $S={\mathbb{C}}[x_{0},x_{1},\dots,x_{n}]$. First of all, let us recall some notions related to this correspondence.
Given a form $f$ of degree $d$, the minimum number of linear forms $l_{i}$ needed to write $f$ as a sum of $d$-th powers is the so-called (Waring) rank of $f$ and we denote it by $\operatorname{rank}(f)$. Note that one can define $\operatorname{rank}([f])$ by $\operatorname{rank}(f)$, because this rank is invariant under non-zero scaling. The (Waring) border rank is given by the same notion in the limiting sense. In other words, if there is a family $\{f_{\epsilon}\mid\epsilon>0\}$ of polynomials with constant rank $r$ and $\lim_{\epsilon\to 0}f_{\epsilon}=f$, then we say that $f$ has border rank at most $r$. The minimum such $r$ is called the border rank of $f$ and we denote it again by $\underline{\mathrm{rank}}(f)$. Note that by definition $\sigma_{k}(v_{d}({\mathbb{P}}V))$ is the variety of homogeneous polynomials $f$ of degree $d$ with border rank $\underline{\mathrm{rank}}(f)\leq k$.
@@
$g\cdot f=g(\partial_{0},\partial_{1},\ldots,\partial_{n})f\in S_{a},$
@@
$\mathbf{X}^{I}=\frac{1}{i_{0}!\cdots i_{n}!}x_{0}^{i_{0}}\cdots x_{n}^{i_{n}}$ $\quad\textrm{and}\quad\mathbf{Y}^{J}=y_{0}^{j_{0}}\cdots y_{n}^{j_{n}},$
@@
($\ &b_{I,J}&\\ &&\end{array}\right)\quad\textrm{with $b_{I,J}=b_{I+J}$},$
in the bases defined above. We call this ‘the $(d-a,a)$-symmetric flattening (or catalecticant) matrix’ of $f$. It is easy to see that the transpose $\phi_{d-a,a}(f)^{T}$ is equal to $\phi_{a,d-a}(f)$.
It is obvious that if $f$ has (border) rank 1, then any symmetric flattening $\phi_{d-a,a}(f)$ has rank 1. By subadditivity of matrix rank, we also know that $\operatorname{rank}~{}\phi_{d-a,a}(f)\leq r$ if $\underline{\mathrm{rank}}(f)\leq r$. So, we obtain a set of defining equations coming from $(k+1)$-minors of the matrix $\phi_{d-a,a}(f)$ for $\sigma_{k}(v_{d}({\mathbb{P}}V))$. For some small values of $k$, it is known that these minors are sufficient to cut the variety $\sigma_{k}(v_{d}({\mathbb{P}}V))$ scheme-theoretically (see [LO13, Theorem 3.2.1]).
Let us recall some more basic terms and facts. Let $Z\subset{\mathbb{P}}W$ be a (reduced and irreducible) variety and $\hat{Z}$ be its affine cone in $W$. Consider a (closed) point $\hat{p}\in\hat{Z}$ and say $p$ the corresponding point in ${\mathbb{P}}W$. We denote the affine tangent space to $Z$ at $p$ in $W$ by $\hat{T}_{p}Z$ and we define the (affine) conormal space to $Z$ at $p$, $\hat{N}^{\ast}_{p}Z$ as the annihilator $(\hat{T}_{p}Z)^{\perp}\subset W^{\ast}$. Since $\dim\hat{N}^{\ast}_{p}Z+\dim\hat{T}_{p}Z=\dim W$ and $\dim Z\leq\dim\hat{T}_{p}Z-1$, we get that
(21) $\dim\hat{N}^{\ast}_{p}Z\leq\operatorname{codim}(Z,{\mathbb{P}}W)$
and the equality holds if and only if $Z$ is smooth at $p$. This conormal space is quite useful to study the (embedded) tangent space $\mathbb{T}_{p}Z$.
For any given form $f\in S^{d}V$, we call $\partial\in T_{t}$ apolar to $f$ if the differentiation $\partial\cdot f$ gives zero (i.e., $\partial\in\ker\phi_{t,d-t}(f)$). And we define the apolar ideal $f^{\perp}\subset T$ as
$f^{\perp}=\{\partial\in T~{}|~{}\partial\cdot f=0\}.$
It is straightforward to see that $f^{\perp}$ is indeed an ideal of $T$. Moreover, it is well-known that the quotient ring $T_{f}=T/f^{\perp}$ is an Artinian Gorenstein algebra with socle degree $d$ (see e.g. [IK99, Chapter 1]). In terms of this apolar ideal, we have a useful description of (a part of) conormal space as follows:
@@
Suppose that $f\in S^{d}V$ corresponds to a (closed) point $[f]$ of $\sigma_{k}(v_{d}({\mathbb{P}}V))\setminus\sigma_{k-1}(v_{d}({\mathbb{P}}V))$. Then, for any $a$ with $1\leq a\leq{\lfloor\frac{d+1}{2}\rfloor}$ with $\operatorname{rank}\phi_{d-a,a}(f)=k$ we have
@@
Let $X\subset{\mathbb{P}}W$ be any variety. For any linear embedding $W\hookrightarrow A\otimes B$ and the induced embedding $X\subset{\mathbb{P}}W\hookrightarrow{\mathbb{P}}(A\otimes B)$, it is well-known that for any $[f]\in\sigma_{k}(X)\subset{\mathbb{P}}(A\otimes B)$, considering $\sigma_{k}(X)$ as a subvariety of ${\mathbb{P}}(A\otimes B)$, we have
$\hat{N}^{\ast}_{[f]}\sigma_{k}(X)\supseteq\ker(f)\otimes\operatorname{im}(f)^{\perp}=\hat{N}^{\ast}_{[f]}\sigma_{p}(Seg({\mathbb{P}}A\times{\mathbb{P}}B))$
in $A^{\ast}\otimes B^{\ast}$ provided that $X\subseteq\sigma_{p}(Seg({\mathbb{P}}A\times{\mathbb{P}}B))$, $X\nsubseteq\sigma_{p-1}(Seg({\mathbb{P}}A\times{\mathbb{P}}B))$ and $f$ has rank $k\cdot p$ as a linear map in $\operatorname{Hom}(A^{\ast},B)$ (see e.g. [LO13, section 2.5]). Here, $Seg({\mathbb{P}}A\times{\mathbb{P}}B)$ means the Segre variety in ${\mathbb{P}}(A\otimes B)$.
@@
The assertion is immediate when we apply this fact to a partial polarization $S^{d}V\hookrightarrow S^{a}V\otimes S^{d-a}V$, because $X=v_{d}({\mathbb{P}}V)$ is contained in $Seg({\mathbb{P}}S^{a}V\times{\mathbb{P}}S^{d-a}V)\subset{\mathbb{P}}(S^{a}V\otimes S^{d-a}V)$ (i.e., $p=1$ case) and $\operatorname{rank}\phi_{d-a,a}(f)=k$, $\ker\phi_{d-a,a}(f)=(f^{\perp})_{d-a}$, $\operatorname{im}(\phi_{d-a,a}(f))^{\perp}=(f^{\perp})_{a}$. ∎
@@ 3.2. Proof of Theorem 1
(i) Let $f$ be any form belonging to $\sigma_{k}(v_{d}({\mathbb{P}}^{1}))\setminus\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$. Set $X=v_{d}({\mathbb{P}}^{n})\subset{\mathbb{P}}^{N}$, the Veronese variety. Consider $f$ as a polynomial in ${\mathbb{C}}[x_{0},x_{1}]$ as in §3.1. Then, by [IK99, Theorem 1.44] we know that $T/f^{\perp}$ is an Artinian Gorenstein algebra with socle degree $d$ and that $f^{\perp}$ is a complete intersection of two homogeneous polynomials $F,G$ of each degree $a$ and $b$ $(a\leq b)$ with $a+b=d+2$ as an ideal of ${\mathbb{C}}[y_{0},y_{1}]$, where the Hilbert function of $T/f^{\perp}$ is
(22) $(1,2,\ldots,a-1,a,\ldots,a,a-1,\ldots,2,1).$
We claim that $\operatorname{rank}~{}\phi_{k,d-k}(f)=k$ (i.e., $a=k$). If $a<k$, then by the shape (22) we see that $\operatorname{rank}~{}\phi_{t,d-t}(f)<k$ for all $t$. In particular, all $k$-minors of $\phi_{t,d-t}(f)$ vanish for any $t$. Since the $k$-minors of catalecticant $\phi_{t,d-t}$ for each $k-1\leq t\leq d-(k-1)$ give the ideal of $\sigma_{k-1}(v_{d}({\mathbb{P}}^{1}))$ (e.g. [Ibid., Theorem 1.45]), this implies $f\in\sigma_{k-1}(v_{d}({\mathbb{P}}^{1}))\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$, which is a contradiction.
@@
Now, let us show that $\sigma_{k}(X)$ is smooth at $f$ by computing the dimension of conormal space. In general, by (21) we have
(23) (${n+d\choose d}-kn-k\geq\dim_{\mathbb{C}}\hat{N}^{\ast}_{[f]}\sigma_{k}(X),$
where the left hand side is given by the expected codimension of the $k$-th secant variety. By Proposition 23, we also have
(24) $\dim_{\mathbb{C}}\hat{N}^{\ast}_{[f]}\sigma_{k}(X)\geq\dim_{\mathbb{C}}(f^{\perp})_{k}\cdot(f^{\perp})_{d-k}\,.$
@@
Since $k\leq\frac{d+1}{2}$ by the assumption, note that $d-k\geq k$ unless $d$ is odd and $k=\frac{d+1}{2}$, where $d-k=\frac{d-1}{2}<k$.
a) If $d$ is odd and $k=\frac{d+1}{2}$, then we have
@@
b) When $d$ is odd and $k<\frac{d+1}{2}$ or $d$ is even, we have $k\leq d-k$ and
@@
(ii) First note that $\dim\sigma_{k}(v_{d}({\mathbb{P}}^{1}))=\min\left\{2k-1,d\right\}$ and the incidence $I$ has dimension $12k-$ 1. In the case $d\leq 2k-2$, each fiber of $p:I\rightarrow{\mathbb{P}}^{d}$ is of dimension $\geq 1$, so for a general $a\in\sigma_{k}(v_{d}({\mathbb{P}}^{1}))$, it holds $q_{i}(p^{-1}(a))=v_{d}({\mathbb{P}}^{1})$ for some $i$ with $1\leq i\leq k$ in the incidence (9) in §2.2.
Now, let $n\geq 3,k=3$ or $n\geq 2,k\geq 4$ and $d=2k-2$. Suppose $\sigma_{k}(v_{d}({\mathbb{P}}^{1}))\not\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$. Then a general point $a\in\sigma_{k}(v_{d}({\mathbb{P}}^{1}))={\mathbb{P}}^{d}$ is a smooth point of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$. Since $q_{i}(p^{-1}(a))=v_{d}({\mathbb{P}}^{1})$ for some $i$, it follows from Lemma 12 that for $M=\mathbb{T}_{a}\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$, we have the inclusion $\mathbb{T}_{x}(v_{d}({\mathbb{P}}^{n}))\subset M$ for a general $x\in v_{d}({\mathbb{P}}^{1})$, and then the inclusion holds for any $x\in v_{d}({\mathbb{P}}^{1})$. This is because, for the Gauss map $\gamma:v_{d}({\mathbb{P}}^{n})\rightarrow{\mathbb{G}}(n,{\mathbb{P}}^{N})$ sending $\gamma(z)=\mathbb{T}_{z}(v_{d}({\mathbb{P}}^{n}))$ (it is a morphism since $v_{d}({\mathbb{P}}^{n})$ is smooth), considering the closed set $G_{M}=\left\{W\in{\mathbb{G}}(n,{\mathbb{P}}^{N})\mid W\subset M\right\}$, we have $\gamma(U)\subset G_{M}$ for a certain non-empty open subset $U\subset v_{d}({\mathbb{P}}^{1})$, and then $\gamma(v_{d}({\mathbb{P}}^{1}))\subset G_{M}$. Therefore,
(25) $\langle\,\bigcup_{x\in v_{d}({\mathbb{P}}^{1})}\mathbb{T}_{x}(v_{d}({\mathbb{P}}^{n}))\,\rangle\subset\mathbb{T}_{a}\sigma_{k}(v_{d}({\mathbb{P}}^{n})).$
Taking $m=1$, $\Lambda=\emptyset$, and $A={\mathbb{P}}^{1}$ in Proposition 22, the number (19), a lower bound for dimension of left hand side of (25), is equal to $dn$. Thus we have
$(2k-2)n=dn\leq k(n+1)-1\;\big{(}\!=\dim\mathbb{T}_{a}\sigma_{k}(v_{d}({\mathbb{P}}^{n}))\big{)},$